OpenStudy (anonymous):

how does 1/20 + 1/15 equal 8.57

OpenStudy (lgbasallote):

60 is the LCD.. (4 + 3)/60 = 7/60 = 0.1166 should be that...

OpenStudy (anonymous):

well... I think they're not equal \[\LARGE \frac{1}{20}+\frac{1}{15}=8\frac{57}{100}\] \[\LARGE \frac{3}{60}+\frac{4}{60}=\frac{800}{100}+\frac{57}{100}\] \[\LARGE \frac{7}{60}=\frac{857}{100}\] they're not equal !

OpenStudy (anonymous):

http://pages.uoregon.edu/dparks/206/resistorvoltage/

OpenStudy (anonymous):

can u guys try and figure out what this guy did pls

OpenStudy (anonymous):

which task ?

OpenStudy (anonymous):

and that's not math, that's physic !

OpenStudy (lgbasallote):

it's 1/R lol..you take the reciprocal :p

OpenStudy (lgbasallote):

1/R = 7/60 R = 60/7 R = what you said

OpenStudy (anonymous):

1/Ra=1/20+1/15 Ra=8.57Ω this one?? :B

OpenStudy (anonymous):

yea that one

OpenStudy (anonymous):

\[\LARGE \frac{1}{Ra}=\frac{1}{20}+\frac{1}{15}\] \[\LARGE \frac{1}{Ra}=\frac{7}{60}\] \[\LARGE 60=7\cdot Ra \quad \quad \Longrightarrow Ra=\frac{60}{7}\] Ra=8.571428...

OpenStudy (anonymous):

lol i see

OpenStudy (anonymous):

ok